# Mind the memory: Consistent time reversal removes artefactual scaling of energy dissipation rate and provides more accurate and reliable thermodynamic inference

### v3

@@ Abstract
It has been proposed that an observed inverse power-law dependence of the Markovian estimate for the steady-state dissipation rate on the coarse-graining scale in self-similar networks reflects a scale-dependent energy dissipation. By explicit examples, it is demonstrated here that there are in general no relations between such an apparent power-law dependence and the actual dissipation on different length scales. We construct fractal networks with a single dissipative scale and networks with a true inverse energy-dissipation cascade, and show that they display the same scaling behavior. Moreover, we show that a self-similar network structure does not imply an inverse power-law scaling but may be mistaken for one in practice. When no dissipative cycles become hidden by the coarse graining, any scale dependence of the dissipation estimate vanishes if the memory is correctly accounted for in the time-reversal operation. A $k$-th order estimator is derived and necessary and sufficient conditions are proved for a guaranteed lower bound on dissipation. These higher-order estimators saturated in the order are proved to provide sharper lower bounds on dissipation and their scale dependence signifies hidden dissipative cycles. It is shown that estimators not saturated in the order may erroneously overestimate the microscopic dissipation. Our results underscore the still underappreciated importance of correctly accounting for memory in analyzing coarse observations.
Measurements in complex systems usually have finite resolution [1, 2, 3, 4] or do not resolve at all relevant degrees of freedom—they just probe some low-dimensional projections of the dynamics, where many microscopic states become “lumped” onto the same observable ”state”. Typical examples are macroscopic observables in condensed matter systems, such as magnetization [5], dielectric response [6], and diverse order parameters [7, 8, 9], as well as other observables (e.g. molecular extensions or FRET efficiencies and lifetimes) in single-molecule [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and particle-tracking [20, 21] experiments.
@@
$\dot{S}[\Gamma_{t}]=k_{\rm B}\lim_{t\to\infty}\frac{1}{t}\left\langle\ln\frac{\operatorname{\mathbb{P}}\left[\Gamma_{t}\right]}{\operatorname{\mathbb{P}}\left[\theta\Gamma_{t}\right]}\right\rangle,$ (1)
[…]
@@
When a coarse-grained version $\hat{X}_{t}$ of the process is observed, where several microscopic states are lumped into the same observable state, provided a physically consistent time reversal $\theta$ [26, 32, 45, 39], the definition in Eq. (1) with lumped paths $\hat{\Gamma}_{t}=(\hat{X}_{\tau})_{0\leq\tau\leq t}$ still holds, whereas Eq. (2) does not. The reason is that $\hat{X}_{t}$ displays memory (due to coarse-graining) which must be taken explicitly into account [46, 32, 33, 34, 39, 47]. If one simply replaces the rate $L_{ji}$ between microscopic states in the Markovian result (2) with the effective rates between (mesoscopic) lumped states averaged over $P^{\rm s}$, i.e. implicitly assuming the existence of a (potentially unphysical [33]) infinite timescale separation between dynamics within and between lumped states, one finds [28] ${\dot{S}_{\rm M}[\hat{\Gamma}_{t}]\leq\dot{S}_{\rm M}[\Gamma_{t}]}$. However, this inequality does not imply that a dependence of $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]$ on the properties of the lumped state space carries any physical meaning since $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]$ is not the dissipation rate of the coarse-grained process, ${\dot{S}_{\rm M}[\hat{\Gamma}_{t}]\neq\dot{S}[\hat{\Gamma}_{t}]}$. Thermodynamic interpretations in absence of an infinite timescale separation of $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]$ (see, e.g. [48, 49, 50]) may lead to misconceptions, as we explicitly show below.
Figure 1: Model systems: (a) Cycle graph with transition rates $\omega$ and $\kappa$ in counter-clockwise ($+$) and clockwise (--) direction, respectively. Lumps $L_{1},L_{2},\ldots,L_{n/\lambda}$ are highlighted in blue.(b) Tree of depth $d=5$ with lumps of depth $l=1$ (shaded in blue) and depth $l=2$ (circled in dashed red). Transition rates down $+$ and up -- (cyclic with respect to the thick edge) are set ${\propto\omega}$ and ${\propto\kappa}$, respectively, chosen such that the cumulative transition rate between each level is ${\omega-\kappa}$. Self-similarity becomes obvious upon noticing that each tree consists of multiple smaller trees. Inset: Four fundamental cycles passing through the highlighted subgraph on levels $U_{3}$, $U_{4}$ and corresponding transition rates. (c) Self-similar Sierpinski-type graph with lumps of size $\lambda=12$ indicated by the respective vertex coloring. A stationary current $\omega$ runs through the outer side of each polygon, whereby the size of said polygons depends on the respective recursive depth. The depicted graph has recursion depth $5$ and $n=768$ vertices. Inset: Element of recursion-depth $2$ with all irreversible currents indicated by arrows. (d) Brusselator depicted as a grid graph as in [48]. A portion of the total $n=450\times 450$ states is shown with lumps of size $\lambda=2\times 2$ and $\lambda=4\times 4$ shaded in blue and circled in dashed red, respectively.
@@
To this end, recent works [48, 69, 49, 50] investigated how the energy dissipation rate in self-similar non-equilibrium reaction networks depends on the coarse-graining scale $\lambda\equiv n/n_{s}$ (i.e. the total number of microstates $n$ relative to the number of mesoscopic states $n_{s}$ in a lumped system) in absence of a timescale separation and observed a power-law dependence of the Markovian estimate $\dot{S}_{\rm M}[\hat{\Gamma}^{s}_{t}]\propto\lambda^{-\alpha}$ [48]. These findings were rationalized in terms of an “inverse cascade” of energy dissipation on different scales. In the specific context of active flows generated by the microtubule-kinesin motor system [70], this amounts to the claim that the most energy is spent to generate and maintain the flow at smaller length scales and only a tiny amount at large length scales [48]. While the underlying calculations in Refs. [48, 69, 49, 50] are technically sound within the inconsistent Markovian approximation (the authors of [69] were aware of this issue in contrast to [48]), their interpretation [48, 49] is physically inconsistent, as we show in this Letter.
Here we construct explicit examples invalidating such hypothesized implications between a power-law dependence of the Markovian dissipation estimates $\dot{S}_{\rm M}$ and the actual energy dissipated on different scales. In particular, we present fractal networks dissipating (i) only on a single (i.e., largest) length-scale and (ii) networks dissipating mostly on the smallest length scale and less on larger scales (i.e., networks with an “inverse energy dissipation cascade” [48]) whereby both display the same scaling of $\dot{S}_{\rm M}$. A self-similar network structure is shown to not even imply a power-law scaling of $\dot{S}_{\rm M}$, therefore there is generally no relation between the power-law scaling of $\dot{S}_{\rm M}$ and the energy dissipated on a given scale. Moreover, the observed (apparent) power-law exponent changes upon progressively accounting for memory in the estimate for $\dot{S}$. The apparent power-law dependence of $\dot{S}_{\rm M}$ thus has no relation to the actual energy dissipated on a given scale and may simply be an artifact of an inconsistent time reversal. Extrapolations to the microscopic scale $\lambda=1$ are shown to be generally non-informative. Most importantly, when no dissipative cycles become hidden by the coarse graining, any scale dependence vanishes entirely once the memory is accounted for exactly when estimating $\dot{S}$. Higher-order dissipation estimates are proved to be reliable indicators of hidden dissipative cycles and, if saturated in the order, provide guaranteed sharper lower bounds. These results emphasize the importance of properly accounting for memory effects in analyzing measurements from coarse observations.
@@
where $T$ is the average waiting time per state in the observed trajectories, $\theta\hat{\gamma}_{k}$ is the corresponding reversed set of $k+1$ states, and the upper indices refer to the discrete state sequence  [^1] That is, the sequence of states of the embedded Markov chain.. If $\hat{\Gamma}_{t}$ is a $n$-th order semi-Markov process then $\dot{S}^{\rm est}_{k}=\dot{S}^{\rm est}_{n},\,\forall k\geq n$ (see proof in [73]), and if $\hat{\Gamma}_{t}=\Gamma_{t}$ is a Markov process we have $\dot{S}^{\rm est}_{1}=\dot{S}$. Notably, we have that if the observed semi-Markov process $\hat{\Gamma}$ is of order $n$, then $\dot{S}^{\rm est}_{k}$ for $k\geq n$ provides a lower bound to the microscopic entropy production rate
@@
implying, alongside $\dot{S}^{\rm est}_{k}=\dot{S}^{\rm est}_{n}\;\forall k\geq n$, that the estimator will reach a plateau upon increasing $k$ once reaching the correct order $n$. Furthermore, $\dot{S}^{\rm est}_{n}$ (but not $\dot{S}^{\rm est}_{k<n}$) always provides a lower bound on $\dot{S}(\Gamma)$ (see proof in [73]; see also [74] for results in discrete time). Crucially, no statement can be made about the monotonicity of $\dot{S}^{\rm est}_{k}$ as a function of $k$ for $k<n$; the inequality (4) may be violated (see counterexample in Appendix A). Inferring the exact $\dot{S}$ from projected observations is generally not possible because of dissipative cycles hidden in lumps.
@@
As in [48], we decimate the state space in $n/\lambda$ coarse-grained observable states (we assume $n=0\,{\rm mod}\,\lambda$), with “Markovian” forward and backward transition rates according to [48] given by $\Omega(\lambda)=\omega/\lambda$ and ${\rm K}(\lambda)=\kappa/\lambda$, respectively. We obtain a single dissipative cycle with homogeneous steady-state current $J=(\lambda/n)[\Omega(\lambda)-{\rm K}(\lambda)]=(\omega-\kappa)/n$ and affinity $A=(n/\lambda)\ln(\omega/\kappa)$, yielding the Markov estimate
@@
Example 2: Tree graph.— We now consider a perfectly self-similar network with a single dissipative scale (i.e. the system size) but multiple (equivalent) dissipative cycles each with a stationary current $J_{c}$ (see Fig. 1b). In this binary tree of depth $d$, every edge is bidirectional with rate $\propto\omega$ in the “down” ($+$) and $\propto\kappa$ in the “up” (--) direction, respectively, such that the cumulative transition rate between each level is ${\omega-\kappa}$. This way, the rates along adjacent edges in consecutive levels in both directions $\pm$ are halved (see Inset in  1b). The detailed mathematical construction is described in the SI. The exact entropy production of the system in the cycle basis $\mathcal{C}$ with $2^{d}$ basis cycles reads (see proof in SM)
@@
where $\Phi_{\pm|\pm}$ is the conditional probability for a transition between two lumps (at given coarseness) in the $\pm$ direction given that the preceding step occurred in the $\pm$ direction. According to the last equality, consistent time reversal (and hence correct thermodynamics) removes the artefactual scaling behavior. The analytical results are confirmed by simulations (see 2b) and generalized to the more complex tree-diamond graph in Appendix C.
@@
Figure 3: Entropy-production estimates for different coarse-graining scales: (a) $\dot{S}^{\rm est}$ for Sierpinski-type graph for recursion depth 7 (${n=12288}$ vertices). We find $\dot{S}_{\rm M}\propto\lambda^{-1}$ while higher-order estimators $\dot{S}^{\rm est}_{k\geq 2}$ yield a power law with ${\alpha=0.93}$, corroborating numerically that the lumped process is a $2^{\rm nd}$ order semi-Markov process. (b) $\dot{S}^{\rm est}$ for the Brusselator with $n=202500$ vertices. The virtual power law for the Markov estimate ${\dot{S}_{1}^{\rm est}\propto\lambda^{-0.52}}$ is faster decaying than $\dot{S}^{\rm est}_{k\geq 2}\propto\lambda^{-0.31}$ obtained by accounting for memory. For simulation details see [73].
@@
Example 4: Brusselator.—Finally, let us consider the Brusselator model [76, 77, 78] (see Fig. 1d) as in [48], here with 202500 microscopic states. We perform the coarse graining as was done in [48], two levels of coarse graining are highlighted in Fig. 1d. We estimate $\dot{S}$, $\dot{S}_{\rm M}$, and $\dot{S}^{\rm est}_{k\geq 2}$ from 25 simulated stationary trajectories each having $10^{9}$ steps (skipping $10^{9}$ initial steps). Using Eq. (3) with ${k=1}$ we infer $\dot{S}_{\rm M}$ that agrees with the results in [48] (red triangles in Fig. 3b). In particular, we reproduce the apparent power-law scaling $\dot{S}_{\rm M}\propto\lambda^{-\alpha}$ with the same exponent as in [48], i.e. $\alpha\approx 0.52$. Coincidentally, it is of the same order as the virtual power-law of the tree and tree-diamond with $\alpha\approx 0.3$ (see Appendix C), which in reality is not a power law at all (see Eq. (9)).
@@
We now account for memory in the time-reversal via $\dot{S}^{\rm est}_{k\geq 2}$. The lumped process is apparently ${2^{\rm nd}}$ or ${3^{\rm rd}}$ order semi-Markov as ${\dot{S}^{\rm est}_{2,3}=\dot{S}^{\rm est}_{4}}$. The fact that ${\dot{S}^{\rm est}_{k\geq 2}<\dot{S}}$ is a consequence of dissipative cycles that become hidden by the coarse graining. Higher-order estimates also display an apparent power-law scaling albeit with a smaller exponent ${\dot{S}^{\rm est}_{\geq 2}\propto\lambda^{-0.31}}$ for considered $\omega,\kappa$. However, according to previous counterexamples, we must conclude that these power laws in fact have no implications about energy dissipation on different length scales.
Interpreting scalings under consistent time reversal.—An apparent power-law dependence of $\dot{S}_{\rm M}$ on the coarse-graining scale $\lambda$ generally has no relation with the microscopic dissipation mechanisms. However, meaningful information about the underlying dynamics may still be inferred from coarse-grained data under consistent time reversal. More specifically, if we can find an order $n$ such that $\dot{S}^{\rm est}_{n+k}=\dot{S}^{\rm est}_{n}$ for any $k\geq 1$. If for such $n,k$ we find that $\dot{S}^{\rm est}_{n+k}\propto\lambda^{-\alpha}$ for some $\alpha>0$, one concludes that the coarse graining hides dissipative cycles (this is not true for a scaling of $\dot{S}_{\rm M}$). If in addition, we have information supporting a self-similar network structure, we may also assume that most energy is dissipated on the smallest and less energy on larger scales.
@@
Conclusion.—Several explicit examples that invalidate the recently proposed relation between an apparent power-law dependence of Markovian estimates for the energy dissipation rate $\dot{S}_{\rm M}$ from coarse-grained observations and the actual dissipation on different length scales were presented and analyzed. We constructed fractal networks with a single dissipative scale and self-similar networks with an “inverse energy dissipation cascade” [48], where most energy is dissipated on the smallest and less energy on larger scales. Both systems exhibit the same scaling behavior of the energy dissipation at different scales. The apparent scaling exponent gradually reduces upon accounting for memory in the dissipation estimates. The hypothesized power-law scaling [48] has already been considered in experiment but has not been validated [79]; our results may explain why this may not be possible. When no dissipative cycles become hidden by the coarse graining, any scale dependence vanishes as soon as the memory is accounted for exactly in the time-reversal operation and we can infer the microscopic dissipation. Therefore, an inverse power-law dependence of the dissipation rate generally has no implications for scale-dependent energy dissipation and may simply be an artifact of inconsistent time reversal in presence of memory. However, it may still be possible to infer meaningful information about the microscopic dynamics, e.g. the presence of hidden dissipative cycles, via higher-order dissipation estimators established here. These incorporate a consistent time reversal in the presence of memory and, if (and only if) saturated in the order, yield experimentally accessible sharper and guaranteed lower bounds on dissipation, further highlighting the importance and benefits of accounting for memory in analyzing data from coarse observations.
@@
Figure 4: (a) 6-state Markov process coarse-grained to process of order $n>3$. The order is at least 3 because the next transition out of $L_{3}$ depends on whether we arrived from lump $L_{1}$ or $L_{4}$ two state-changes before. Since multiple jumps between $L_{2}\leftrightarrow L_{3}$ may have occurred before, the order is in fact higher, but the influence of these back-and-forth jumps vanishes exponentially. (b) Entropy production rate estimated with $\dot{S}^{\rm est}_{k}$ as a function of $k$ from trajectories of different length. While undersampling effects occur for higher orders (red and blue lines do not agree for $k\geq 8$), the overshoot for lower orders is not an artifact of undersampling. This example explicitly shows that second-order estimators are not necessarily a lower bound [50].

### v4

# Consistent time reversal and reliable and accurate inference in the presence of memory

@@ Abstract
Thermodynamic inference from coarse observations remains a key challenge. Memory, in particular correlations between consecutively observed mesostates, blur signatures of irreversibility and must be accounted for in defining physical time-reversal, which remains an open problem. We derive an experimentally accessible $k$-th order estimator for the entropy production rate. Using novel measure-theoretic techniques we prove necessary and sufficient conditions for guaranteed lower bounds on the dissipation even in the strongly non-Markovian setting. The proof reveals that estimators saturated in the order unravel the duration of memory which needs to be considered in defining physically consistent time-reversal. We show that Markovian estimators in absence of a time-scale separation lead to artifacts, which convey no physical meaning. Similarly, estimators not saturated in the order may overestimate the dissipation. The necessity of correctly accounting for memory in thermodynamic inference from strongly non-Markovian observations underscores the still underappreciated challenges and intricacies in defining and understanding irreversibility in presence of memory. Our results will hopefully stimulate experiments systematically considering thermodynamic inference on multiple scales consistently accounting for memory.
Measurements in complex systems usually have finite resolution [1, 2, 3, 4] or do not resolve at all relevant degrees of freedom—they just probe some low-dimensional projections of the dynamics, where many microscopic states become “lumped” onto the same observable “state”. Examples are macroscopic observables in condensed matter systems, such as magnetization [5], dielectric response [6], and diverse order parameters [7, 8, 9], as well as other observables (e.g. molecular extensions or FRET efficiencies and lifetimes) in neural [10], single-molecule [11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and particle-tracking [21, 22] experiments.
@@
$\dot{S}[\Gamma_{t}]=k_{\rm B}\lim_{t\to\infty}\frac{1}{t}\left\langle\ln\frac{\mathrm{d}\operatorname{\mathbb{P}}\left[\Gamma_{t}\right]}{\mathrm{d}\operatorname{\mathbb{P}}\left[\theta\Gamma_{t}\right]}\right\rangle,$ (1)
where $\langle\cdot\rangle$ denotes the average over ${\rm d}{\mathbb{P}}[\Gamma_{t}]$ and $k_{\rm B}$ is Boltzmann’s constant. When $x_{t}$ is an overdamped Markov dynamics, we simply have $\theta\Gamma_{t}=(x_{t-\tau})_{0\leq\tau\leq t}$. Specifically, for an ergodic Markov process on a discrete state space $V$ with transition rates $L_{ji}$ from state $i$ to $j$ and stationary probabilities $P^{\rm s}$, Eq. (1) becomes [43, 44, 45]
@@
When a coarse-grained version $\hat{X}_{t}$ of the process is observed, where several microscopic states are lumped into the same observable state, provided a physically consistent time reversal $\theta$ [27, 33, 46, 40], the definition in Eq. (1) with lumped paths $\hat{\Gamma}_{t}=(\hat{X}_{\tau})_{0\leq\tau\leq t}$ may still hold, whereas Eq. (2) does not. The reason is that $\hat{X}_{t}$ displays memory due to coarse-graining, i.e., correlations between consecutively observed mesostates, which must be taken explicitly into account in the time-reversal [47, 33, 34, 35, 40, 48]. Determining the precise manner in which memory must enter the consistent time reversal remains elusive. If one simply replaces the rate $L_{ji}$ between microscopic states in the Markovian result (2) with the effective rates between (mesoscopic) lumped states averaged over $P^{\rm s}$, i.e. implicitly assuming the existence of a (potentially unphysical [34]) infinite timescale separation between dynamics within and between lumped states one finds [29] $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]\leq\dot{S}_{\rm M}[\Gamma_{t}]$}. However, this inequality does not imply that a dependence of $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]$ on the properties of the lumped state space carries any physical meaning since $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]$ is not the dissipation rate of the coarse-grained process, $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]\neq\dot{S}[\hat{\Gamma}_{t}]$}. In absence of an infinite timescale separation thermodynamic interpretations of $\dot{S}_{\rm M}[\hat{\Gamma}_{t}]$ (see, e.g. [49, 50]) may lead to misconceptions and guaranteed higher-order estimators $\dot{S}^{\rm est}_{k}$ are required, as we explicitly show below. A time-scale separation on some scale does not imply a separation on coarser scales. Under time-scale separation on all scales, $\dot{S}_{\rm M}$ does not display any scale dependence, see Appendix D.
Figure 1: Coarse graining (left) and estimators (right) for $u=v=2$: $\dot{S}^{\mathrm{est}}_{k}$ allows for good inference of $\dot{S}$ already for $k=2$ and saturates in the order for $k\geq 3$.
@@
Physically inconsistent time reversal ignoring memory leads to potential misinterpretation of results [49]. More importantly, memory implies correlation between consecutively visited mesostates but the extend and manner in which these need to be accounted for when defining time reversal symmetry, and in turn when estimating the entropy production rate, remains a grand challenge.
We prove here with measure-theoretic techniques that the only relevant transition probabilities that enter the estimation of $\dot{S}$ are those accounting for the actual extent of memory (but not those accounting for shorter or longer memory). This means that saturation in the order is an inherent quantifier for the extent of memory which enters the definition of physical time-reversal. We construct an experimentally accessible estimator of $\dot{S}$ for strongly non-Markovian observations (see Example in Fig. 1) and prove that it yields guaranteed sharp lower bounds.
We demonstrate the necessity of the new framework by invalidating the hypothesized [49] implications between a power-law dependence of the Markovian dissipation estimates $\dot{S}_{\rm M}$ and the actual energy dissipated on different scales. In particular, we present fractal networks dissipating (i) only on a single (i.e., largest) length-scale and (ii) networks dissipating mostly on the smallest length scale and less on larger scales (i.e., networks with an “inverse energy dissipation cascade” [49]) whereby both display the same scaling of $\dot{S}_{\rm M}$. A self-similar network structure is shown to not even imply a power-law scaling of $\dot{S}_{\rm M}$, therefore there is generally no relation between the power-law scaling of $\dot{S}_{\rm M}$ and the energy dissipated on a given scale. Using our new estimator, the observed (apparent) power-law exponent changes upon progressively accounting for memory in the estimate for $\dot{S}$. The apparent power-law dependence of $\dot{S}_{\rm M}$ thus has no relation to the actual energy dissipated on a given scale and may simply be an artifact of an inconsistent time reversal, i.e. neglecting correlations between consecutively visited mesostates due to memory. Extrapolations to the microscopic scale $\lambda=1$ are shown to be generally non-informative. Most importantly, when no dissipative cycles become hidden by the coarse graining, any scale dependence vanishes entirely once the memory is accounted for exactly when estimating $\dot{S}$. Furthermore, by introducing a measure-theoretic framework, our higher-order dissipation estimates are proved to be reliable indicators of hidden dissipative cycles and, if saturated in the order which ensures consistent time-reversal, provide guaranteed sharper lower bounds. These results emphasize the importance of properly accounting for memory effects in analyzing measurements from coarse observations.
@@
where $T$ is the average waiting time per state in the observed trajectories, $\theta\hat{\gamma}_{k}$ is the corresponding reversed set of $k+1$ states, and the upper indices refer to the discrete state sequence  [^1] That is, the sequence of states of the embedded Markov chain.. We prove in [72] that $\dot{S}^{\rm uncond}_{k}=\sum_{i=1}^{k}\dot{S}^{\rm est}_{i}$ and highlight in Appendix A pitfalls of using $\dot{S}^{\rm uncond}_{k}$. If $\hat{\Gamma}_{t}$ is a $n$-th order semi-Markov process then $\dot{S}^{\rm est}_{k}=\dot{S}^{\rm est}_{n},\,\forall k\geq n$ (see proof in [72]), and if $\hat{\Gamma}_{t}=\Gamma_{t}$ is a Markov process we have $\dot{S}^{\rm est}_{1}=\dot{S}$. Notably, if the observed semi-Markov process $\hat{\Gamma}$ is of order $n$, then $\dot{S}^{\rm est}_{k}$ for $k\geq n$ provides a lower bound on the microscopic $\dot{S}$
@@
implying, alongside $\dot{S}^{\rm est}_{k}=\dot{S}^{\rm est}_{n}\;\forall k\geq n$, that the estimator will reach a plateau upon increasing $k$ once reaching the correct order $n$; at this order (and not before) consistent time-reversal is obtained, correctly accounting for correlations between consecutive mesostates. Furthermore, $\dot{S}^{\rm est}_{n}$ (but not $\dot{S}^{\rm est}_{k<n}$) always provides a lower bound on $\dot{S}(\Gamma)$ (see our proof in [72]; this has been stated, but not proven, for discrete time in [73]). As shown in Appendix E, in continuous time mesoscopic paths may be a projection of uncountably many microscopic paths, creating the need for a novel measure-theoretic technique that we develop here. Our proof includes the discrete time result. Crucially, no statement can be made about the monotonicity of $\dot{S}^{\rm est}_{k}$ as a function of $k$ for $k<n$; the inequality (4) may be violated (see example in Appendix A). Inferring the exact $\dot{S}$ from projected observations is generally not possible because of dissipative cycles hidden in lumps.
Sharpening inference at strong memory.—First, we provide a system in Fig. 1 that illustrates the benefits of accounting for long-range memory in thermodynamic inference: We observe that accounting for progressively longer memory $k$ yields a better estimate of $\dot{S}$ via $\dot{S}^{\mathrm{est}}_{k}$, until we reach saturation for $k\geq 3$. Therefore, physically consistent time-reversal requires accounting for $k=3$ past states. Note the substantial improvement over the Markovian estimate $\dot{S}_{\rm M}=\dot{S}^{\rm est}_{1}$ [49]. Further, while $\dot{S}^{\rm uncond}_{2}/2$ [51] provides a lower bound (this estimator was originally developed for the broader setting of blurred transitions [74]), we recover with $\dot{S}^{\mathrm{est}}_{2}$ in Fig. 1 $75\%$ more at the same statistical need, and get saturation for $k\geq 3$ (upto statistical fluctuations). Further parameters are considered in Appendix F. An efficient GPU-implementation of the estimator is provided. We now demonstrate the necessity of accounting for memory in time reversal.
Figure 2: Model systems: (a) Cycle graph with transition rates $\omega$ and $\kappa$ in counter-clockwise (++) and clockwise (-) direction, respectively. Lumps $L_{1},L_{2},\ldots,L_{n/\lambda}$ are highlighted in blue.(b) Tree of depth $d=5$ with lumps of depth $l=1$ (shaded in blue) and depth $l=2$ (circled in dashed red). Transition rates down ++ and up - (cyclic with respect to the thick edge) are set $\propto\omega$} and $\propto\kappa$}, respectively, chosen such that the cumulative transition rate between each level is $\omega-\kappa$}. Self-similarity becomes obvious upon noticing that each tree consists of multiple smaller trees. Inset: Four fundamental cycles passing through the highlighted subgraph on levels $U_{3}$, $U_{4}$ and corresponding transition rates. (c) Self-similar Sierpinski-type graph with lumps of size $\lambda=12$ indicated by the respective vertex coloring. A stationary current $\omega$ runs through the outer side of each polygon, whereby the size of said polygons depends on the respective recursive depth. The depicted graph has recursion depth $5$ and $n=768$ vertices. Inset:Furthermore, a corresponding sequence of increasing jumptimes in $[0,t]$ must exist.lement of recursion-depth 22 with all irreversible currents indicated by arrows. (d) Brusselator depicted as a grid graph as in [49]. A portion of the total $n=450\times 450$ states is shown with lumps of size $\lambda=2\times 2$ (shaded in blue) and $\lambda=4\times 4$ (circled in dashed red).
@@
As in [49], we decimate the state space in $n/\lambda$ observable mesostates (we assume $n=0\,{\rm mod}\,\lambda$), with “Markovian” forward and backward rates according to [49] given by $\Omega(\lambda)=\omega/\lambda$ and ${\rm K}(\lambda)=\kappa/\lambda$, respectively. We obtain a single dissipative cycle with homogeneous steady-state current $J=(\lambda/n)[\Omega(\lambda)-{\rm K}(\lambda)]=(\omega-\kappa)/n$ and affinity $A=(n/\lambda)\ln(\omega/\kappa)$, yielding the Markov estimate
@@
Example 2: Tree graph.— We now consider a perfectly self-similar network with a single dissipative scale (i.e. the system size) but multiple (equivalent) dissipative cycles each with a stationary current $J_{c}$ (see Fig. 2b). In this binary tree of depth $d$, every edge is bidirectional with rate $\propto\omega$ in the “down” (++) and $\propto\kappa$ in the “up” (-) direction, respectively, such that the cumulative transition rate between each level is $\omega-\kappa$}. This way, the rates along adjacent edges in consecutive levels in both directions $\pm$ are halved (see Inset in  2b). The detailed mathematical construction is described in the SI. The exact entropy production of the system in the cycle basis $\mathcal{C}$ with $2^{d}$ basis cycles reads (see proof in SM)
@@
where $\Phi_{\pm|\pm}$ is the conditional probability for a transition between two lumps (at given coarseness) in the $\pm$ direction given that the preceding step occurred in the $\pm$ direction. According to the last equality, consistent time reversal (and hence correct thermodynamics) removes the artefactual scaling behavior. The analytical results are confirmed by simulations (see 3b) and generalized to the more complex tree-diamond graph in Appendix C.
@@
Figure 4: Entropy-production estimates for different coarse-graining scales: (a) $\dot{S}^{\rm est}$ for Sierpinski-type graph for recursion depth 7 (n=12288{n=12288} vertices). We find $\dot{S}_{\rm M}\propto\lambda^{-1}$ while higher-order estimators $\dot{S}^{\rm est}_{k\geq 2}$ yield a power law with $\alpha=0.93$}, corroborating numerically that the lumped process is a $2^{\rm nd}$ order semi-Markov process. (b) $\dot{S}^{\rm est}$ for the Brusselator with $n=202500$ vertices. The virtual power law for the Markov estimate $\dot{S}_{1}^{\rm est}\propto\lambda^{-0.52}$} is faster decaying than $\dot{S}^{\rm est}_{k\geq 2}\propto\lambda^{-0.31}$ obtained by accounting for memory. For details, see [72].
@@
Example 4: Brusselator.—Finally, let us consider the Brusselator model [76, 77, 78] (see Fig. 2d) as in [49], here with 202500 microscopic states. We perform the coarse graining as was done in [49], two levels of coarse graining are highlighted in Fig. 2d. We estimate $\dot{S}$, $\dot{S}_{\rm M}$, and $\dot{S}^{\rm est}_{k\geq 2}$ from 25 simulated stationary trajectories each having $10^{9}$ steps (skipping $10^{9}$ initial steps). Using Eq. (3) with k=1{k=1} we infer $\dot{S}_{\rm M}$ that agrees with the results in [49] (red triangles in Fig. 4b). In particular, we reproduce the apparent power-law scaling $\dot{S}_{\rm M}\propto\lambda^{-\alpha}$ with the same exponent as in [49], i.e. $\alpha\approx 0.52$. Coincidentally, it is of the same order as the virtual power-law of the tree and tree-diamond with $\alpha\approx 0.3$ (see Appendix C), which in reality is not a power law at all (see Eq. (9)).
@@
We now account for memory in the time-reversal via $\dot{S}^{\rm est}_{k\geq 2}$. The lumped process is apparently $2^{\rm nd}$} or $3^{\rm rd}$} order semi-Markov as $\dot{S}^{\rm est}_{2,3}=\dot{S}^{\rm est}_{4}$}. The fact that $\dot{S}^{\rm est}_{k\geq 2}<\dot{S}$} is a consequence of dissipative cycles that become hidden by the coarse graining. Higher-order estimates also display an apparent power-law scaling albeit with a smaller exponent $\dot{S}^{\rm est}_{\geq 2}\propto\lambda^{-0.31}$} for considered $\omega,\kappa$. However, according to previous examples, we must conclude that these power laws in fact have no implications about energy dissipation on different length scales (see Section IX. [72]).
@@
Conclusion.—A novel, experimentally accessible dissipation estimator for coarse-grained observations in continuous-time was introduced and proved that it yields a guaranteed lower-bound. The estimator is applicable to strongly non-Markovian observations (i.e. beyond weakly Markovian [79, 80] and correlation-time expansions [81, 82]). For the first time, the estimator provides insight into the extent of memory that needs to be considered when defining physically consistent time-reversal. The necessity of correctly accounting for memory in time-reversal was demonstrated by invalidating a recently proposed relation between an apparent power-law dependence of Markovian estimates for the energy dissipation rate $\dot{S}_{\rm M}$ from coarse-grained observations and the actual dissipation on different length scales were presented and analyzed. We constructed fractal networks with a single dissipative scale and self-similar networks with an “inverse energy dissipation cascade” [49], where most energy is dissipated on the smallest and less energy on larger scales. Both systems exhibit the same scaling behavior of the energy dissipation at different scales. The apparent scaling exponent gradually reduces upon accounting for memory in the dissipation estimates and only if saturated in the order, provides information on hidden dissipative cycles (see [72]); scaling laws ignoring memory reflect a mixture of memory and hidden cycles. The hypothesized power-law scaling [49] has already been considered in experiment but a power law has not been validated [83];  we hope that our novel framework stimulates experiments systematically considering multiple scales. When no dissipative cycles become hidden by the coarse graining, any scale dependence vanishes as soon as the memory is accounted for exactly in the time-reversal operation and we can infer the microscopic dissipation. Therefore, an inverse power-law dependence of the dissipation rate generally has no implications for scale-dependent energy dissipation and may simply be an artifact of inconsistent time reversal in presence of memory. However, it may still be possible to infer meaningful information about the microscopic dynamics, e.g. the presence of hidden dissipative cycles, via higher-order dissipation estimators established here. These incorporate a consistent time reversal in the presence of memory and, if (and only if) saturated in the order, yield experimentally accessible [84, 85], feasible [86, 87], sharper and guaranteed lower bounds on dissipation, further highlighting the importance and benefits of accounting for memory in analyzing data from coarse observations.
@@
Figure 5: (a) 6-state Markov process coarse-grained to process of order $n>3$. The order is at least 3 because the next transition out of $L_{3}$ depends on whether we arrived from lump $L_{1}$ or $L_{4}$ two state-changes before. Since multiple jumps between $L_{2}\leftrightarrow L_{3}$ may have occurred before, the order is in fact higher, but the influence of these back-and-forth jumps vanishes exponentially. (b) Entropy production rate estimated with $\dot{S}^{\rm est}_{k}$ as a function of $k$ from trajectories of different length. While undersampling effects occur for higher orders (red and blue lines do not agree for $k\geq 8$), the overshoot for lower orders is not an artifact of undersampling. This example explicitly shows that second-order estimators are not necessarily a lower bound. Notably, the numerical algorithm for determining all estimators up to order 10 completes within a few seconds (red line) and minutes (blue line) on a normal desktop computer; for details see [72].